OpenStudy (anonymous):

use the quadratic formula to find any x-intercepts on the graph of the equation. y=5x^2+6x-6

OpenStudy (anonymous):

so far I have \[\frac{ -6\pm \sqrt{156} }{ 10 }\]

OpenStudy (mertsj):

156 = 2x2x3x13 Simplify the radical using that information.

OpenStudy (anonymous):

thats where I am getting stuck..

OpenStudy (mertsj):

\[\sqrt{156}=\sqrt{2\times2 \times3\times13}=\sqrt{4\times3\times13}=\sqrt{4}\sqrt{3}\sqrt{13}\]

OpenStudy (mertsj):

\[\sqrt{4}\sqrt{3}\sqrt{13}=2\sqrt{3}\sqrt{13}=2\sqrt{39}\]

OpenStudy (anonymous):

Now I understand! thank you so much for you help and patience. @Mertsj

OpenStudy (mertsj):

yw